Jeff is twice as old as Sherry. Their combined age is 72. How old is each person?
Answer by ikleyn(53762)

.

Let x be the Sherry's age.


Then the Jeff's age is 2x years, according to the condition.


The equation for the total age is


    x + 2x = 72.


Simplify and solve


    3x = 72

     x = 72/3 = 24.


ANSWER.  Sherry is 24 years old.  Jeff is 2*24 = 48 years.

Solved.